# PyTorch Frame: A Modular Framework for Multi-Modal Tabular Learning

### v1

@@ 1 Introduction
However, GBDT has notable limitations. First, GBDT models are primarily focused on numerical and categorical features and cannot effectively handle raw multi-modal features, such as texts, sequences, images, and embeddings. Second, their end-to-end integration with downstream deep learning models, such as Graph Neural Networks (GNNs), is highly non-trivial since GBDT models are neither differentiable nor producing embeddings (Ivanov & Prokhorenkova, 2021). As such, GBDT falls short on complex applications, such as prediction over modern relational databases (Fey et al., 2023).
@@ 2 PyTorch Frame
PyTorch Frame111github.com/pyg-team/pytorch-frame provides a unified framework for efficient deep learning over tabular data $\mathbf{T}=[(v_{1},\ldots,v_{C})]^{N}_{n=1}$, which holds data across $C$ columns for every of its $N$ rows. We denote $T[i,j]$ as the raw value of column $j$ in row $i$. We also use standard NumPy notations (Harris et al., 2020), such as $T[:,j]$, $T[i,:]$, $T[[i_{1},...,i_{k}],:]$, and $T[:,[j_{1},...,j_{k}]]$.
@@ 2.1 Materialization
text_embedded. Similar to text_tokenized, each cell of $\mathbf{T}[:,\phi(s)]$ consists of a piece of text. Different from text_tokenized, we use external text embedding models (Neelakantan et al., 2022; Reimers & Gurevych, 2019) to pre-compute text vectors. Concretely, each single column $\mathbf{T}[:,j],j\in\phi(s)$ is pre-encoded by a column-specific text model into an embedding tensor of shape $[N,1,D_{j}]$, where $D_{j}$ can be different for different $j$. To handle multiple text columns simultaneously, PyTorch Frame introduces its own MultiEmbeddingTensor layout, where each column of $\mathbf{T}[:,\phi(s)]$ is pre-embedded into $D_{j}$-dimensional vectors, which are stacked to produce MultiEmbeddingTensor $\mathbf{F}_{s}$ of shape $[N,|\phi(s)|,\mathbf{D}]$. We also support image_embedded, where each cell contains the path of the image data.
@@ 2.2 Encoding
Column Embeddings. After feature normalization, the encoders embed $\mathbf{F}_{s}$, representing $|\phi(s)|$ columns, into $F$-dimensional column embeddings $\mathbf{X}_{j}$ of shape $[N,|\phi(s)|,F]$. Different modeling choices are possible. For example, numerical columns can either be transformed either via a linear layer or can be first converted into piecewise linear or periodic representations (Gorishniy et al., 2022) before a linear layer. For categorical columns, one can transform them via shallow embeddings learnable for each category. The text_tokenized columns can be transformed into embeddings via language models that take the sequences of tokens as input.
@@ 2.3 Column-wise Interaction
Many existing works can be cast under this framework. For example, Gorishniy et al. (2021) applied a permutation-invariant Transformer (Vaswani et al., 2017) to model column interactions, while Huang et al. (2020) used a Transformer with positional column encoding. (Chen et al., 2023a) also followed a Transformer architecture except that it sorts the features by mutual information and used diagonal attention in the Transformer block. Chen et al. (2023b) used cross attention between column embeddings and learnable prompt embeddings to model column interactions.
@@ 3 Integration
Integration with PyTorch Geometric. We have so far discussed single-table tabular learning, but many practical applications involve data stored in a relational format (Codd, 1970), where tabular data is connected with each other via primary-foreign key relationships. Combining tabular deep learning with Graph Neural Networks (GNNs) has proven to be promising to handle such relational datasets (Fey et al., 2023). PyTorch Frame integrates natively with PyTorch Geometric (PyG) (Fey & Lenssen, 2019), a popular PyTorch library for GNNs. PyTorch Frame enhances PyG by learning embedding vectors of nodes and edges with complex multi-modal features. The node and edge embeddings are subsequently fed as input to GNNs by PyG. Crucially, tabular deep learning models by PyTorch Frame and GNNs by PyG can be jointly trained to optimize for downstream task performance.
@@ 4.1 Handling single-table data
Text model Tabular Model Text Stype ROCAUC Time ROCAUC Time ROCAUC Time
ResNet embedded 0.934 7.3s 0.883 36.1s 0.753 27.8s
RoBERTa FTTransformer embedded 0.936 19.7s 0.882 100.8s 0.747 77.4s
(Embedding) Trompt embedded 0.958 18.8s 0.885 480.4s 0.756 581.9s
LightGBM embedded 0.954 15.5s 0.865 571.1s 0.767 1931.9s
OpenAI ResNet embedded 0.923 10.4s 0.945 56.5s 0.807 107.1s
FTTransformer embedded 0.911 23.6s 0.945 337.4s 0.807 168.9s
Trompt embedded 0.976 40.8s 0.947 4285.1s 0.810 538.0s
LightGBM embedded 0.966 131.0s 0.926 1732.9s 0.809 1924.3s
RoBERTa ResNet tokenized 0.979 5.5h 0.970 $>$1 day 0.786 $>$1 day
(Finetune) FTTransformer tokenized 0.960 5.5h 0.968 $>$1 day 0.775 $>$1 day
Best single model from Shi et al. (2021) 0.967 - 0.967 - 0.794 -

### v2

@@ 1 Introduction
However, GBDT has notable limitations. First, GBDT models are primarily focused on numerical and categorical features and cannot effectively handle raw multi-modal features, such as texts, sequences, images, and embeddings. Second, their end-to-end integration with downstream deep learning models, such as Graph Neural Networks (GNNs), is highly non-trivial since GBDT models are neither differentiable nor producing embeddings (Ivanov & Prokhorenkova, 2021). As such, GBDT falls short on complex applications, such as prediction over modern relational databases Fey et al. (2023).
@@
2 Related Work
Our framework follows the modular encoder-combiner-decoder framework (Molino et al., 2019), while being explicit about modeling multi-layer column interactions notable in modern deep tabular models (Chen et al., 2023a; Gorishniy et al., 2021; Chen et al., 2023b; Huang et al., 2020). Our framework is also related to PyTorch Tabular (Joseph, 2021), an open-source tabular learning framework built on top PyTorch. While PyTorch Tabular has primarily focused on supporting existing tabular models, our PyTorch Frame offers enhanced flexibility for exploring and building novel tabular learning approaches while still providing access to established models. PyTorch Frame further distinguishes itself through support for a wider variety of column modalities and streamlined integration with LLMs.
@@ 3 PyTorch Frame
PyTorch Frame  [^1] https://github.com/pyg-team/pytorch-frame provides a unified framework for efficient deep learning over tabular data $\mathbf{T}=[(v_{1},\ldots,v_{C})]^{N}_{n=1}$, which holds data across $C$ columns for every of its $N$ rows. We denote $T[i,j]$ as the raw value of column $j$ in row $i$. We also use standard NumPy notations (Harris et al., 2020), such as $T[:,j]$, $T[i,:]$, $T[[i_{1},...,i_{k}],:]$, and $T[:,[j_{1},...,j_{k}]]$.
@@ 3.1 Materialization
text_embedded. Similar to text_tokenized, each cell of $\mathbf{T}[:,\phi(s)]$ consists of a piece of text. Different from text_tokenized, we use external text embedding models (Neelakantan et al., 2022; Reimers & Gurevych, 2019) to pre-compute text vectors. Concretely, each single column $\mathbf{T}[:,j],j\in\phi(s)$ is pre-encoded by a column-specific text model into an embedding tensor of shape $[N,1,D_{j}]$, where $D_{j}$ can be different for different $j$. To handle multiple text columns simultaneously, PyTorch Frame introduces its own MultiEmbeddingTensor layout, where each column of $\mathbf{T}[:,\phi(s)]$ is pre-embedded into $D_{j}$-dimensional vectors, which are stacked to produce MultiEmbeddingTensor $\mathbf{F}_{s}$ of shape $[N,|\phi(s)|,\mathbf{D}]$. We also support image_embedded in the similar way, whereas each cell contains the path of the image data.
@@ 3.2 Encoding
Column Embeddings. After feature normalization, the encoders embed $\mathbf{F}_{s}$, representing $|\phi(s)|$ columns, into $F$-dimensional column embeddings $\mathbf{X}_{j}$ of shape $[N,|\phi(s)|,F]$. Different modeling choices are possible. For example, numerical columns can either be transformed either via a linear layer or can be first converted into piecewise linear or periodic representations Gorishniy et al. (2022) before a linear layer. For categorical columns, one can transform them via shallow embeddings learnable for each category. The text_tokenized columns can be transformed into embeddings via language models that take the sequences of tokens as input.
@@ 3.3 Column-wise Interaction
Many existing works can be cast under this framework. For example, Gorishniy et al. (2021) applied a permutation-invariant Transformer (Vaswani et al., 2017) to model column interactions, while Huang et al. (2020) used a Transformer with positional column encoding. Chen et al. (2023a) also followed a Transformer architecture except that it sorts the features by mutual information and used diagonal attention in the Transformer block. Chen et al. (2023b) used cross attention between column embeddings and learnable prompt embeddings to model column interactions.
@@ 4 Integration
Integration with PyTorch Geometric. We have so far discussed single-table tabular learning, but many practical applications involve data stored in a relational format (Codd, 1970), where tabular data is connected with each other via primary-foreign key relationships. Combining tabular deep learning with Graph Neural Networks (GNNs) has proven to be promising to handle such relational datasets Fey et al. (2023). PyTorch Frame integrates natively with PyTorch Geometric (PyG) Fey & Lenssen (2019), a popular PyTorch library for GNNs. PyTorch Frame enhances PyG by learning embedding vectors of nodes and edges with complex multi-modal features. The node and edge embeddings are subsequently fed as input to GNNs by PyG. Crucially, tabular deep learning models by PyTorch Frame and GNNs by PyG can be jointly trained to optimize for downstream task performance.
@@ 5.1 Handling single-table data
Text Model Tabular Model ROC-AUC Time ROC-AUC Time ROC-AUC Time
RoBERTa ResNet 0.934 7.3s 0.883 36.1s 0.753 27.8s
FTTransformer 0.936 19.7s 0.882 100.8s 0.747 77.4s
Trompt 0.958 18.8s 0.885 480.4s 0.756 581.9s
LightGBM 0.954 15.5s 0.865 571.1s 0.767 1931.9s
OpenAI ResNet 0.923 10.4s 0.945 56.5s 0.807 107.1s
FTTransformer 0.911 23.6s 0.945 337.4s 0.807 168.9s
Trompt 0.976 40.8s 0.947 4285.1s 0.810 538.0s
LightGBM 0.966 131.0s 0.926 1732.9s 0.809 1924.3s
RoBERTa ResNet 0.979 5.5h 0.970 $>$1 day 0.786 $>$1 day
FTTransformer 0.960 5.5h 0.968 $>$1 day 0.775 $>$1 day
Best single model (Shi et al., 2021) 0.967 - 0.967 - 0.794 -